# ColorSense: A Study on Color Vision in Machine Visual Recognition

### v2

@@ Abstract
Color vision is essential for human visual perception, but its impact on machine perception is still underexplored. There has been an intensified demand for understanding its role in machine perception for safety-critical tasks such as assistive driving and surgery but lacking suitable datasets. To fill this gap, we curate multipurpose datasets ColorSense, by collecting 110,000 non-trivial human annotations of foreground and background color labels from popular visual recognition benchmarks. To investigate the impact of color vision on machine perception, we assign each image a color discrimination level based on its dominant foreground and background colors and use it to study the impact of color vision on machine perception. We validate the use of our datasets by demonstrating that the level of color discrimination has a dominating effect on the performance of mainstream machine perception models. Specifically, we examine the perception ability of machine vision by considering key factors such as model architecture, training objective, model size, training data, and task complexity. Furthermore, to investigate how color and environmental factors affect the robustness of visual recognition in machine perception, we integrate our ColorSense datasets with image corruptions and perform a more comprehensive visual perception evaluation. We jointly analyze the impact of color vision and image corruption on machine perception. Our findings suggest that object recognition tasks such as classification and localization are susceptible to color vision bias, especially for high-stakes cases such as vehicle classes, and advanced mitigation techniques such as data augmentation and so on only give marginal improvement. Our analyses highlight the need for new approaches toward the performance evaluation of machine perception models in real-world applications. Lastly, we present various potential applications of ColorSense such as studying spurious correlations.
@@ III-B Defining Color Discrimination Groups
We follow the Web Content Accessibility Guidelines (WCAG) [1] and consult experts from the School of Fine Arts to assign all possible color pairs into three color discrimination (CD) groups, {Hard, Medium, Esay}. The Hard groups consist of images with the same foreground and background colors. The Easy group includes color pairs with a score larger than the WCAG Minimum Contrast Success Criterion 1.4.3 [1] and they also correspond to the top one-third scores ranked among all pairs; the rest goes into Medium. Tab. V (in Appendix) presents the statistics for the number of color pairs in each CD group and the number of test images in each group for ColorSense datasets. We discuss the label quality and provide details of the color pairs for each color discrimination group in the Appendix.
@@ IV-A Fundamental Questions
The relationship between model size and robustness is an intriguing topic of research in the field of deep learning. Hendrycks et al. [28] showed that larger models tend to be more robust to added noises. This finding is particularly interesting since larger DNNs do resemble the number of neurons in the human brain more closely. Furthermore, human vision is known to be robust to color variations, which raises the question of whether larger DNNs are also more robust to color variations. Exploring the relationship between model size, robustness, and color variation could lead to useful insights into the functioning of DNNs and their potential applications. We compare (1) architectures of the same family and (2) models of similar sizes but different architectures to answer this question in our problem setup.
@@ IV-C Advanced Analyses and Ablations
Real-world applications of DNNs often also involve environmental factors. A benefit of ColorSense is the easy incorporation with OOD datasets like ImageNet-C to task models with added environmental factors and noises. We observe a consistent trend: color vision plays an important role in all types of corruption. In Fig. 6, we show three models, Convnext_s, Swin_b, SwinV2_s of similar accuracies on ImageNet, 83.61, 83.58, and 83.71, respectively (less than 0.15 difference) for comparisons. Even though the generalization performances are almost the same, for each corruption type, the color vision performance is different. For instance, Swin & SwinV2’s color discrimination capabilities are better than Convnext_s for all corruptions but jpeg compression, and SwinV2_s is slightly better than Swin_b on digital corruption such as pixelate but worse on noise corruptions. These findings have implications and insights for decisions such as which model we should choose? If we expect noisy environments for some custom application, we should choose SwinV2_s model among the three.
@@
TABLE IV: Object localization accuracy. The color vision effect exists in another visual recognition task like object localization, not only in classification.
@@
Inspired by [30], we hypothesize that color vision bias arise from certain frequency components of the images. We conduct frequency analysis and split the frequency components with $r=$ [8, 16, 32]. We split each image into high and low frequency components (HFC/LFC), and test Resnet50 and ViT_L on the two components respectively. Results for each CD group are summarized in Tab. LABEL:tab:frequency. We observe that when presented with only the LFC of images, models consistently perform worse in Hard group, and surprisingly, on HFC both models show better performance for Hard group. Since LFC results resemble what we observe for the original images, we hypothesize that the color vision bias is predominately driven by the LFC in the images. Furthermore, as we already observed lower performance in harder groups on the grayscale experiments (Fig. 9) and considering LFC usually contains most of the color information [22], we conjecture that the color vision bias may be relevant to other low-level features that models pick up, beyond mere colors. Consequently, mitigating this bias is non-trivial, as also supported by experiments from previous sections (Tab. I & Fig. 4 & Fig. 5), which also substantiate the merit of our dataset as this phenomenon cannot be explored before.
@@
Accompanying each image is the corresponding attention map, we observe that the model may not be able to attend to the right object areas in whole, or sometimes the object is blended in the background. Thus, we believe that the models’ misclassifications stem primarily from the small color difference between foreground and background elements, which appear to impede the models’ ability to effectively discern and isolate salient features. This phenomenon underscores the ongoing challenges in developing robust computer vision systems capable of matching human perceptual acuity across diverse visual scenarios.
@@ V Other Utilities: Object Localization and Spurious Correlation
We apply the same analysis procedure in Sec. IV-B to object localization [8]. We report top-1 localization accuracy with three different IoU thresholds (30, 50, 70). We summarize the results on ColorSenese-ImageNet in Tab. IV. We observe similar performance differences as in classification task, where the model tends to perform better as the CD group becomes easier. And this trend happens across all three different IoU thresholds. This again bolsters our finding of the color vision effect in machine vision.
@@
VI Discussion
@@
We believe that in the era of foundation models with various scaling laws, our dataset and analysis framework are valuable and can push the boundaries of benchmarking.
VII Potential Broader Impact

### v3

@@ Abstract
Color vision is essential for human visual perception, but its impact on machine perception is still underexplored. There has been an intensified demand for understanding its role in machine perception for safety-critical tasks such as assistive driving and surgery but lacking suitable datasets. To fill this gap, we curate multipurpose datasets ColorSense, by collecting 110,000 non-trivial human annotations of foreground and background color labels from popular visual recognition benchmarks. To investigate the impact of color vision on machine perception, we assign each image a color discrimination level based on its dominant foreground and background colors and use it to study the impact of color vision on machine perception. We validate the use of our datasets by demonstrating that the level of color discrimination has a dominating effect on the performance of mainstream machine perception models. Specifically, we examine the perception ability of machine vision by considering key factors such as model architecture, training objective, model size, training data, and task complexity. Furthermore, to investigate how color and environmental factors affect the robustness of visual recognition in machine perception, we integrate our ColorSense datasets with image corruptions and perform a more comprehensive visual perception evaluation. We jointly analyze the impact of color vision and image corruption on machine perception. Our findings suggest that object recognition tasks such as classification and localization are susceptible to color vision bias, especially for high-stakes cases such as vehicle classes, and advanced mitigation techniques such as data augmentation and so on only give marginal improvement. Our analyses highlight the need for new approaches toward the performance evaluation of machine perception models in real-world applications. Lastly, we present various potential applications of ColorSense such as studying spurious correlations. [^1] Dataset is available at https://github.com/charismaticchiu/ColorSense [^2] This work has been accepted for publication in the IEEE Conference on Secure and Trustworthy Machine Learning (SaTML). The final version will be available on IEEE Xplore.
@@ III-B Defining Color Discrimination Groups
We follow the Web Content Accessibility Guidelines (WCAG) [1] and consult experts from the School of Fine Arts to assign all possible color pairs into three color discrimination (CD) groups, {Hard, Medium, Esay}. The Hard groups consist of images with the same foreground and background colors. The Easy group includes color pairs with a score larger than the WCAG Minimum Contrast Success Criterion 1.4.3 [1] and they also correspond to the top one-third scores ranked among all pairs; the rest goes into Medium. Tab. VI (in Appendix) presents the statistics for the number of color pairs in each CD group and the number of test images in each group for ColorSense datasets. We discuss the label quality and provide details of the color pairs for each color discrimination group in the Appendix.
@@ IV-A Fundamental Questions
The relationship between model size and robustness is an intriguing topic of research in the field of deep learning. Hendrycks et al. [28] showed that larger models tend to be more robust to added noises. Furthermore, human vision is known to be robust to color variations [16, 71, 51, 32, 55, 42, 59, 18, 54, 21], which raises the question of whether larger DNNs are also more robust to color variations. Exploring the relationship between model size, robustness, and color variation could lead to useful insights into the functioning of DNNs and their potential applications. We compare (1) architectures of the same family and (2) models of similar sizes but different architectures to answer this question in our problem setup.
@@ IV-C Advanced Analyses and Ablations
Real-world applications of DNNs often also involve environmental factors. A benefit of ColorSense is the easy incorporation with OOD datasets like ImageNet-C to task models with added environmental factors and noises. We observe a consistent trend: color vision plays an important role in all types of corruption. In Fig. 6, we show three models, Convnext_s, Swin_b, SwinV2_s of similar accuracies (%) on ImageNet, 83.61, 83.58, and 83.71, respectively (less than 0.15 difference) for comparisons. Even though the generalization performances are almost the same, for each corruption type, the color vision performance is different. For instance, Swin & SwinV2’s color discrimination capabilities are better than Convnext_s for all corruptions but jpeg compression, and SwinV2_s is slightly better than Swin_b on digital corruption such as pixelate but worse on noise corruptions. These findings have implications and insights for decisions such as which model we should choose? If we expect noisy environments for some custom application, we should choose SwinV2_s model among the three.
@@
TABLE IV: High & low frequency components accuracy (%). Results reported in Resnet50 / ViT_L format.
LFC Easy Medium Hard $r=8$ 4.04 / 9.64 3.44 / 9.12 3.23 / 8.04 $r=16$ 21.47 / 36.49 19.27 / 34.60 17.81 / 31.39 $r=32$ 51.71 / 64.43 49.01 / 62.52 47.48 / 60.61 HFC Easy Medium Hard $r=8$ 44.97 / 51.08 44.62 / 51.61 45.52 / 52.19 $r=16$ 13.99 / 9.58 14.65 / 10.20 15.00 / 11.23 $r=32$ 2.01 / 1.82 1.85 / 1.64 2.57 / 2.67
TABLE V: Object localization accuracy. The color vision effect exists in another visual recognition task like object localization, not only in classification.
@@
Inspired by [30], we hypothesize that color vision bias arise from certain frequency components of the images. We conduct frequency analysis and split the frequency components with $r=$ [8, 16, 32]. We split each image into high and low frequency components (HFC/LFC), and test Resnet50 and ViT_L on the two components respectively. Results for each CD group are summarized in Tab. IV. We observe that when presented with only the LFC of images, models consistently perform worse in Hard group, and surprisingly, on HFC both models show better performance for Hard group. Since LFC results resemble what we observe for the original images, we hypothesize that the color vision bias is predominately driven by the LFC in the images. Furthermore, as we already observed lower performance in harder groups on the grayscale experiments (Fig. 9) and considering LFC usually contains most of the color information [22], we conjecture that the color vision bias may be relevant to other low-level features that models pick up, beyond mere colors. Consequently, mitigating this bias is non-trivial, as also supported by experiments from previous sections (Tab. I & Fig. 4 & Fig. 5), which also substantiate the merit of our dataset as this phenomenon cannot be explored before.
@@
Accompanying each image is the corresponding attention map, though it may not be the best interpretibility technique, we observe that the model may not be able to attend to the right object areas in whole, or sometimes the object is blended in the background. Thus, we believe that the models’ misclassifications stem primarily from the small color difference between foreground and background elements, which appear to impede the models’ ability to effectively discern and isolate salient features. This phenomenon underscores the ongoing challenges in developing robust computer vision systems capable of matching human perceptual acuity across diverse visual scenarios.
@@
IV-D Practicality of Proposed sCVR & sCR Metrics
We want to highlight the generality of how sCVR was formulated. sCVR is defined to measure the variability of model performance between subgroups while taking accuracy into account. In this work, we have 3 different color discrimination groups per our design; however, in other scenarios, researchers may be interested in a different factor, so they have different subgroups. Therefore, we can design a new metric in a general way, such as sCVR to broadly assess the between-subgroup variability. For instance, we similarly defined an sCR (formula above) for measuring robustness to image corruption (the y-axis in Fig. 10-left). We believe that solely using “total accurac” to evaluate models does not provide sufficient granularity, and a performance variability metric is needed for any study design that involves subgrouping. A simple statistical variance does not suffice due to subgroup sample imbalances. We also designed sCVR along with sCR to account for total accuracy since it is commonly regarded as the primary metric measuring model performance.
To validate the practicality of our metric, we want to highlight that in Fig. 10-left, our finding on model robustness to image corruption is consistent with [27] — size-wise, larger models (represented by larger ticks) have higher sCR (more robust). We believe that sCVR and sCR can serve as general metrics for model robustness.
For both sCVR and sCR, we reference the definition of population standard deviation (divided by w, not w-1), rather than sample standard deviation, since our goal is to quantify a variation measurement using all the groups rather than making a specific statistical inference/estimation.
One can also consider model sizes, but that is out of the scope of our work.
@@ V Other Utilities: Object Localization and Spurious Correlation
We apply the same analysis procedure in Sec. IV-B to object localization [8]. We report top-1 localization accuracy with three different IoU thresholds (30, 50, 70). We summarize the results on ColorSenese-ImageNet in Tab. V. We observe similar performance differences as in classification task, where the model tends to perform better as the CD group becomes easier. And this trend happens across all three different IoU thresholds. This again bolsters our finding of the color vision effect in machine vision.
@@
VI Limitations
We study color vision in the scope of static visual perception, and it would be interesting to extend the analysis to dynamic visual perception such as video data.
Furthermore, following our extensive analyses in Sec. IV-C, to address the color vision phenomenon that our work revealed in modern DL architectures, only advanced training such as data augmentation shows marginal benefit, and it would be a promising future direction to develop a bespoke data augmentation method to enhance the mitigation of color vision in modern DL architectures.
VII Discussion
@@
We believe that in the era of foundation models with various scaling laws, our dataset and analysis framework are valuable and can push the boundaries of benchmarking in visual recognition.
VIII Potential Broader Impact